OpenStudy (anonymous):

what is 5x-3y if =9 for y

OpenStudy (shamil98):

Please rephrase your question.

OpenStudy (petiteme):

if x=9?

OpenStudy (anonymous):

solve 5x-3y=9 for y

OpenStudy (shamil98):

-3y = -5x + 9 y = 5/3x -3

OpenStudy (anonymous):

nooooo i mean 5x-3y=9

OpenStudy (shamil98):

Solve for x or y ? ._.

OpenStudy (anonymous):

idk

OpenStudy (petiteme):

aw. that's a big problem

OpenStudy (anonymous):

it says 5x-3y=9 for y i think feel in y for 9

OpenStudy (anonymous):

soo5x-3x9

OpenStudy (anonymous):

hellooo

OpenStudy (petiteme):

so, what really is the equation?